# Quantum teleportation implies symmetry-protected topological order

### v2

@@ 6.4 Valence-bond states for $k\geq 1$ logical qubits
We now describe another family of resource states suitable for the standard teleportation of $k\geq 1$ logical qubits. In contrast to the generalized cluster states $\lvert C_{k}\rangle$ (77) of Sec. 6.3, the unitary $\mathcal{U}$ that prepares the actual resource state $\lvert\Psi_{t}\rangle$ is straightforwardly related to the unitary that prepares the corresponding bulk SPT state $\lvert V_{k}\rangle$ (80).
Figure 8: The valence-bond resource states for $k=1,2,3$ (top to bottom), where two dots connected by a line denote the Bell state $\lvert\mathrm{Bell}\rangle$. The bulk states in the dotted box realize $\lvert V_{k}\rangle$ (80), which are closely related to the resource state $\lvert\Psi_{t}\rangle$ (36) for teleportation. Performing Bell-basis $XX$ and $ZZ$ measurement on each oval teleports the qubits to the rightmost end points of Bell states, up to Pauli errors.
We refer to this family of states as valence-bond states, due to their resemblance to, e.g., the Affleck-Kennedy-Lieb-Tasaki (AKLT) state Affleck et al. (1987). These states are also closely related to the $\alpha$ chains introduced in Ref. 66. Given a 1D chain with an even number of qubits $N$, we define the $k$-fold valence bond state $\lvert V_{k}\rangle$ (80) as
$\lvert V_{k}\rangle\equiv\bigotimes_{j}\,\lvert\mathrm{Bell}\rangle_{2j,2j+2k-1}\,,~{}~{}$ (80)
where $\lvert\mathrm{Bell}\rangle_{a,b}=(\lvert 00\rangle_{a,b}+\lvert 11\rangle_{a,b})/\sqrt{2}$. These states are depicted in Fig. 8 for $k=1,2,3$, where the dotted box encloses the bulk of the chain. The states $\lvert V_{k}\rangle$ (80) are symmetric under the group $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ generated by
$g^{x}_{p}$ $=\prod\limits_{j}\,X_{2\,k\,j+2\,p}\,X_{2\,k\,j+2\,p+1}$ (81a)
$g^{z}_{p}$ $=\prod\limits_{j}\,Z_{2\,k\,j+2\,p}\,Z_{2\,k\,j+2\,p+1}$ (81b)
for $p\in[1,k]$. The corresponding string order parameters (63) act as one of the two generators (6.4) in the the bulk.
We now describe the valence-bond teleportation protocol. Given $k\geq 1$ logical qubits and a target distance $L_{*}$ (13), we choose the minimum number of measurement regions $m\geq 1$ such that $L=\left(2\,m+1\right)k+1\geq L_{*}$. These protocols involve at most nearest-neighbor gates (with $v=1$), and are optimal in that they use the minimum required depth $T=k+1$ (48), so that $L=T+2\,m\,(T-1)$ saturates the teleportation bound (47).
@@
$\lvert\Psi_{0}\rangle$ $=\prod\limits_{n=1}^{k}\lvert\psi_{n}\rangle_{2n-1}\otimes\lvert 0\rangle_{2n}\otimes\prod\limits_{j=2k+1}^{N}\lvert 0\rangle_{j}\,,~{}~{}$ (82)
where $\lvert\psi_{n}\rangle$ is the $n$th logical state on site $i_{n}=2n-1$. All bulk qubits $j\in[2k+1,2(m+1)k]$ are prepared in the initial state $\lvert 0\rangle$, though the sites $j=i_{n}+1$ and $j=N-n$ for $n\in[1,k]$ may be prepared in any state.
The resource state $\lvert\Psi_{t}\rangle$ (36) for teleportation is prepared in $T=k+1$ layers of unitary evolution labelled $\mathcal{U}_{p}$ for $p\in[0,k]$. While the valence-bond state $\lvert V_{k}\rangle$ (80) is prepared by the first $k$ layers alone, the final unitary layer $\mathcal{U}_{k}$ merely converts the symmetry generators in Eq. 6.4 into on-site operators, so both states are SPTs).
@@
$\mathcal{U}_{0}$ $=\Big{(}\prod\limits_{n=1}^{k}\,\mathrm{SWAP}_{2n-1,2n}\,\Big{)}\,\Bigg{[}\prod\limits_{j=k+1}^{(m+1)k}\,\mathcal{B}_{2j-1,2j}\,\Bigg{]}\,\mathds{1}\,,$ (83)
@@
$\mathcal{B}_{a,b}\equiv\mathrm{CNOT}\left(a\to b\right)\,H_{a}\,,~{}~{}$ (84)
where $\mathrm{CNOT}\left(a\to b\right)$ flips the $Z$ state of qubit $b$ if qubit $a$ is in the $Z$ state $\lvert 1\rangle$, and acts trivially otherwise, so that $\mathcal{B}_{a,b}\,\lvert 00\rangle_{a,b}=\lvert\mathrm{Bell}\rangle_{a,b}$. Note that the SWAP gates in $\mathcal{U}_{0}$ (83) can be omitted if we swap the states of sites $12n-$ 1 and $2n$ in $\lvert\Psi_{0}\rangle$ (82), at the cost of decreasing $L$ (13) by one. When preparing the valence-bond state $\lvert V_{k}\rangle$ (80) directly, $\mathcal{U}_{0}$ only contains the bulk term in Eq. 83.
@@
$\mathcal{U}_{p}=\begin{cases}\prod\limits_{j=1}^{(m+1)k}\,\mathrm{SWAP}_{2j+2p-2,2j+2p-1}\,&\,p~{}{\rm odd}\\[2.0pt] \prod\limits_{j=1}^{(m+1)k}\,\mathrm{SWAP}_{2j+2p-3,2j+2p-2}\,&\,p~{}{\rm even}\end{cases}\,,~{}~{}$ (85)
which can be thought of more simply as alternating layers of SWAP gates on even versus odd bonds, where the first layer of SWAP gates acts on even bonds—i.e., the SWAP gates in $\mathcal{U}_{1}$ (85) are staggered by one site with respect to the Bell encoding channels $\mathcal{B}$ in $\mathcal{U}_{0}$ (83).
The product of the untiaries $\mathcal{U}_{k-1}\cdots\mathcal{U}_{0}$ map the state $\lvert\boldsymbol{0}\rangle$ to the valence-bond state $\lvert V_{k}\rangle$ (80). The analogous state $\lvert\Psi_{T-1}\rangle$ for teleportation is captured by Fig. 8 for $k=1,2,3$; however, writing a general expression is inconvenient. The final layer of unitary evolution applies the Bell decoding channel to the qubits in ovals in Fig. 8, and SWAP gates to all qubits on the right end, i.e.,
$\mathcal{U}_{k}=\prod\limits_{j=1}^{km}\mathcal{B}^{\dagger}_{k+2j-1,k+2j}\,\prod\limits_{n=0}^{k-1}\mathrm{SWAP}_{N-2n-1,N-2n},$ (86)
where the SWAP gates on the final $2k$ sites can be omitted at the cost of decreasing $L$ (13) by one. Following the application of all unitaries, one measures $Z_{j}$ on all sites $[k+1,(2m+1)k]$, where every $2k$ consecutive outcomes are determine recovery channels according to $(m_{x,1},m_{z,1},m_{x,2},m_{z,2},\dots,m_{x,k},m_{z,k})$, where $\sigma^{z/x}_{f_{n}}$ is applied to $f_{n}$ if $\sum_{s=1}^{m}m_{x/z,s}\cong 1$ mod $2$.
@@ 7 Outlook
We have shown that all “standard” teleportation protocols that teleport $k$ logical states a distance $L>vT$ (13) are equivalent to preparing a resource state with nontrivial SPT order protected by the symmetry $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ (65) using local unitary evolution for time $T\ll L$, measuring the $2k$ string order parameters (63) in the bulk, and applying Pauli recovery gates (34) based on the outcomes. These protocols are physical in that the logical states $\lvert\psi_{n}\rangle$ are realized on particular physical qubits. We require that no unitaries are applied to a site after it is measured, though we expect that this requirement can be relaxed.
@@
The crucial feature of standard teleportation protocols from the perspective of constraining the resource state is that all outcome-dependent operations are Pauli gates. The recovery group—generated by all physical error-correction operators $\chi$ in $\mathcal{R}$ (18), modulo complex phases—is simply the $k$-qubit Pauli group $\mathbf{P}_{*}{k}/\mathsf{U}\left(1\right)$ for standard teleportation, which forms a projective representation of the symmetry group $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ (65). We expect that our analysis can be extended, mutatis mutandis, to more general recovery operations.
@@
In this sense, our results (namely, Theorem 18) represent a first step toward classifying all quantum teleportation protocols based solely on their recovery (or decoding) channels (18). There are numerous recovery operations to consider beyond the Paulis of standard teleportation: The corresponding group can be continuous and non-Abelian (i.e., $\chi_{g}\,\chi_{h}\neq\omega(g,h)\,\chi_{h}\,\chi_{g}$ for some complex phase $\omega$). We can also condition measurements (or the basis of a measurement) on prior outcomes, and condition operations on the outcomes of measurements in other ways (e.g., nonlinearly). For example, Ref. 67 considers Pauli recovery operations but does not require that they be conditioned on outcome parities, in which case the resource states can be chosen to have SPT order, but the symmetry generators $U(g)$ are no longer “on site” (49). Alternatively, one might consider non-Abelian—and even continuous—recovery groups, whose corresponding resource states are matrix product states Gross and Eisert (2007, 2010) that generally do not have string order Popp et al. (2005). We expect that our methods using the Stinespring formalism will be crucial in analyzing these more exotic protocols and their potential relation to exotic forms of SPT order.
@@ Acknowledgments
We thank Andy Lucas, Charles Stahl, and Chao Yin for useful discussions and collaboration on related work. This work was supported in part by the U.S. Air Force Office of Scientific Research under Grant No. FA9550-21-1-0195 (YH), the Office of Naval Research via Grant No. N00014-23-1-2533 (YH), the Simons Collaboration on Ultra-Quantum Matter via Award No. 651440 (DTS), and the US Department of Energy via Grant No. DE-SC0024324 (AJF).
@@ Appendix A Proof of Proposition 1
For the teleportation conditions 1 to be necessary and sufficient for quantum state transfer (see Def. 1), they must be equivalent to (5) for the initial state $\lvert\Psi_{0}\rangle$ (6) and final state $\lvert\Psi_{T}\rangle$ (7). We prove both directions of implication in turn, using the fact that all operators acting on a qubit can be decomposed onto the Pauli operators $\mathds{1},X,Y,Z$, which are spanned by $X$ and $Z$. The coefficients $\alpha$ and $\beta$ of a logical state $\lvert\psi\rangle$ can be extracted from the expectation values of the Pauli operators (2.2).
@@
We first establish that Eqs. 5, 6, and 7 imply Eq. 1—i.e., Eq. 1 is necessary for state transfer. First, Eq. 7 implies that $X_{f_{n}}$ is the $X$-type logical operator for the $n$th logical qubit acting on the final state $\lvert\Psi_{T}\rangle$ (7), as it interchanges $\alpha_{n}$ and $\beta_{n}$. In the Heisenberg picture, this logical operator evolves under $\mathcal{W}$ according to
$X^{(n)}_{{\rm L}}(T)\,=\,X_{f_{n}}(T)\,=\,\mathcal{W}^{\dagger}\,X_{f_{n}}\,\mathcal{W}\,,~{}~{}$ (A.87)
@@
and now, using the fact that $X^{2}_{i_{n}}=\mathds{1}$ is involutory, we introduce an arbitrary operator $\mathsf{S}_{n,x}$ via
$X^{(n)}_{{\rm L}}(T)\,$ = $=\,\mathds{1}\,\mathcal{W}^{\dagger}X_{f_{n}}\mathcal{W}\,=\,X^{2}_{i_{n}}\,\mathcal{W}^{\dagger}X_{f_{n}}\mathcal{W}$
$=\,X_{i_{n}}\,\left(X_{i_{n}}\,\mathcal{W}^{\dagger}X_{f_{n}}\mathcal{W}\right)\,\equiv\,X_{i_{n}}\,\mathsf{S}_{n,x}\,,~{}~{}$ (A.88)
which implicitly defines the (unitary) operator $\mathsf{S}_{n,x}$ as
$\mathsf{S}_{n,x}\,\equiv\,X_{i_{n}}\,\mathcal{W}^{\dagger}X_{f_{n}}\mathcal{W}\,.$ (A.89)
@@
Leaving the coefficients $\alpha_{j},\beta_{j}$ for $j\neq n$ implicit for convenience, we apply $\mathsf{S}_{n,x}$ (A.89) to the initial state $\lvert\Psi_{0}(\alpha_{n},\beta_{n})\rangle$ (6) and use the first part of Eq. 5 to find
@@
$\mathsf{S}_{n,x}\,\lvert\Psi_{0}(\alpha_{n},\beta_{n})\rangle\,$ $=\,X_{i_{n}}\,\mathcal{W}^{\dagger}X_{f_{n}}\mathcal{W}\,\lvert\Psi_{0}(\alpha_{n},\beta_{n})\rangle$
@@
$=\,X_{i_{n}}\,\mathcal{W}^{\dagger}X_{f_{n}}\lvert\Psi_{T}(\alpha_{n},\beta_{n})\rangle$
@@
$=\,X_{i_{n}}\,\mathcal{W}^{\dagger}\lvert\Psi_{T}(\beta_{n},\alpha_{n})\rangle$
$=\,X_{i_{n}}\lvert\Psi_{0}(\beta_{n},\alpha_{n})\rangle$
$=\,\lvert\Psi_{0}(\alpha_{n},\beta_{n})\rangle\,,~{}~{}$ (A.90)
@@
meaning that $\mathsf{S}_{n,x}$ (A.89) is an element of the initial-state stabilizer group $\operatorname*{stab}\left(*\right){\lvert\Psi_{0}\rangle}$ (3), and is guaranteed to be unitary. Next, applying the Heisenberg-evolved operator $X_{f_{n}}(T)$ (A.87) to the initial state $\lvert\Psi_{0}\rangle$ (6) gives
$X^{(n)}_{{\rm L}}(T)\,\lvert\Psi_{0}(\alpha_{n},\beta_{n})\rangle\,$ $=\,X_{i_{n}}\,\mathsf{S}_{n,x}\,\lvert\Psi_{0}(\alpha_{n},\beta_{n})\rangle$
$=\,X_{i_{n}}\,\lvert\Psi_{0}(\alpha_{n},\beta_{n})\rangle\,,~{}~{}$ (A.91)
@@
which holds for any choice of $\alpha_{n},\beta_{n}$. This implies that $X^{(n)}_{{\rm L}}(T)\,=\,\mathcal{W}^{\dagger}\,X_{f_{n}}\,\mathcal{W}\,$ realizes an $X$-type logical operator for the $n$th logical qubit acting on the initial state $\lvert\Psi_{0}\rangle$ (6)—i.e., it acts as $X_{i_{n}}$ times some stabilizer $\mathsf{S}_{n,x}$ (A.89), where the latter acts trivially on $\lvert\Psi_{0}\rangle$ (6).

### v3

@@ 6.4 Valence-bond states for $k\geq 1$ logical qubits
We now describe another family of resource states suitable for the standard teleportation of $k\geq 1$ logical qubits. In contrast to the generalized cluster states $\lvert C^{\,}_{k}\rangle$ (77) of Sec. 6.3, the unitary $\mathcal{U}$ that prepares the actual resource state $\lvert\Psi^{\,}_{t}\rangle$ is straightforwardly related to the unitary that prepares the corresponding bulk SPT state $\lvert V^{\,}_{k}\rangle$ (80).
Figure 8: The valence-bond resource states for $k=1,2,3$ (top to bottom), where two dots connected by a line denote the Bell state $\lvert\mathrm{Bell}\rangle$. The bulk states in the dotted box realize $\lvert V^{\,}_{k}\rangle$ (80), which are closely related to the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) for teleportation. Performing Bell-basis $XX$ and $ZZ$ measurement on each oval teleports the qubits to the rightmost end points of Bell states, up to Pauli errors.
We refer to this family of states as valence-bond states, due to their resemblance to, e.g., the Affleck-Kennedy-Lieb-Tasaki (AKLT) state [73]. These states are also closely related to the $\alpha$ chains introduced in Ref. 74. Given a 1D chain with an even number of qubits $N$, we define the $k$-fold valence bond state $\lvert V^{\,}_{k}\rangle$ (80) as
$\lvert V^{\vphantom{\prime}}_{k}\rangle\equiv\bigotimes_{j}\,\lvert\mathrm{Bell}\rangle^{\vphantom{\prime}}_{2j,2j+2k-1}\,,~{}~{}$ (80)
where $\lvert\mathrm{Bell}\rangle^{\,}_{a,b}=(\lvert 00\rangle^{\,}_{a,b}+\lvert 11\rangle^{\,}_{a,b})/\sqrt{2}$. These states are depicted in Fig. 8 for $k=1,2,3$, where the dotted box encloses the bulk of the chain. The states $\lvert V^{\,}_{k}\rangle$ (80) are symmetric under the group $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ generated by
$g^{x}_{p}$ $=\prod\limits_{j}\,X^{\vphantom{*}}_{2\,k\,j+2\,p}\,X^{\vphantom{*}}_{2\,k\,j+2\,p+1}$ (81a)
$g^{z}_{p}$ $=\prod\limits_{j}\,Z^{\vphantom{*}}_{2\,k\,j+2\,p}\,Z^{\vphantom{*}}_{2\,k\,j+2\,p+1}$ (81b)
for $p\in[1,k]$. The corresponding string order parameters (63) act as one of the two generators (81) in the the bulk.
We now describe the valence-bond teleportation protocol. Given $k\geq 1$ logical qubits and a target distance $L^{\,}_{*}$ (13), we choose the minimum number of measurement regions $m\geq 1$ such that $L=\left(2\,m+1\right)k+1\geq L^{\,}_{*}$. These protocols involve at most nearest-neighbor gates (with $v=1$), and are optimal in that they use the minimum required depth $T=k+1$ (48), so that $L=T+2\,m\,(T-1)$ saturates the teleportation bound (47).
@@
$\lvert\Psi^{\vphantom{\prime}}_{0}\rangle$ $=\prod\limits_{n=1}^{k}\lvert\psi^{\vphantom{\prime}}_{n}\rangle^{\vphantom{\prime}}_{2n-1}\otimes\lvert 0\rangle^{\vphantom{\prime}}_{2n}\otimes\prod\limits_{j=2k+1}^{N}\lvert 0\rangle^{\vphantom{\prime}}_{j}\,,~{}~{}$ (82)
where $\lvert\psi^{\,}_{n}\rangle$ is the $n$th logical state on site $i_{n}=2n-1$. All bulk qubits $j\in[2k+1,2(m+1)k]$ are prepared in the initial state $\lvert 0\rangle$, though the sites $j=i_{n}+1$ and $j=N-n$ for $n\in[1,k]$ may be prepared in any state.
The resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) for teleportation is prepared in $T=k+1$ layers of unitary evolution labelled $\mathcal{U}^{\,}_{p}$ for $p\in[0,k]$. While the valence-bond state $\lvert V^{\,}_{k}\rangle$ (80) is prepared by the first $k$ layers alone, the final unitary layer $\mathcal{U}^{\,}_{k}$ merely converts the symmetry generators in Eq. 81 into on-site operators, so both states are SPTs).
@@
$\mathcal{U}^{\vphantom{\dagger}}_{0}$ $=\Big{(}\prod\limits_{n=1}^{k}\,\mathrm{SWAP}^{\vphantom{\dagger}}_{2n-1,2n}\,\Big{)}\,\Bigg{[}\prod\limits_{j=k+1}^{(m+1)k}\,\mathcal{B}^{\vphantom{\dagger}}_{2j-1,2j}\,\Bigg{]}\,\mathds{1}\,,$ (83)
@@
$\mathcal{B}^{\vphantom{\dagger}}_{a,b}\equiv\mathrm{CNOT}\left(a\to b\right)\,H^{\vphantom{\dagger}}_{a}\,,~{}~{}$ (84)
where $\mathrm{CNOT}\left(a\to b\right)$ flips the $Z^{\vphantom{*}}$ state of qubit $b$ if qubit $a$ is in the $Z^{\vphantom{*}}$ state $\lvert 1\rangle$, and acts trivially otherwise, so that $\mathcal{B}^{\,}_{a,b}\,\lvert 00\rangle^{\,}_{a,b}=\lvert\mathrm{Bell}\rangle^{\,}_{a,b}$. Note that the SWAP gates in $\mathcal{U}^{\,}_{0}$ (83) can be omitted if we swap the states of sites $12n-$ 1 and $2n$ in $\lvert\Psi^{\,}_{0}\rangle$ (82), at the cost of decreasing $L$ (13) by one. When preparing the valence-bond state $\lvert V^{\,}_{k}\rangle$ (80) directly, $\mathcal{U}^{\,}_{0}$ only contains the bulk term in Eq. 83.
@@
$\mathcal{U}^{\vphantom{\dagger}}_{p}=\begin{cases}\prod\limits_{j=1}^{(m+1)k^{\vphantom{\dagger}}}\,\mathrm{SWAP}^{\vphantom{\dagger}}_{2j+2p-2,2j+2p-1}\,&\,p~{}{\rm odd}\\[2.0pt] \prod\limits_{j=1}^{(m+1)k}\,\mathrm{SWAP}^{\vphantom{\dagger}}_{2j+2p-3,2j+2p-2}\,&\,p~{}{\rm even}\end{cases}\,,~{}~{}$ (85)
which can be thought of more simply as alternating layers of SWAP gates on even versus odd bonds, where the first layer of SWAP gates acts on even bonds—i.e., the SWAP gates in $\mathcal{U}^{\,}_{1}$ (85) are staggered by one site with respect to the Bell encoding channels $\mathcal{B}$ in $\mathcal{U}^{\,}_{0}$ (83).
The product of the untiaries $\mathcal{U}^{\,}_{k-1}\cdots\mathcal{U}^{\,}_{0}$ map the state $\lvert\boldsymbol{0}\rangle$ to the valence-bond state $\lvert V^{\,}_{k}\rangle$ (80). The analogous state $\lvert\Psi^{\,}_{T-1}\rangle$ for teleportation is captured by Fig. 8 for $k=1,2,3$; however, writing a general expression is inconvenient. The final layer of unitary evolution applies the Bell decoding channel to the qubits in ovals in Fig. 8, and SWAP gates to all qubits on the right end, i.e.,
$\mathcal{U}^{\vphantom{\dagger}}_{k}=\prod\limits_{j=1}^{km}\mathcal{B}^{\dagger}_{k+2j-1,k+2j}\,\prod\limits_{n=0}^{k-1}\mathrm{SWAP}^{\vphantom{\dagger}}_{N-2n-1,N-2n},$ (86)
where the SWAP gates on the final $2k$ sites can be omitted at the cost of decreasing $L$ (13) by one. Following the application of all unitaries, one measures $Z^{\vphantom{*}}_{j}$ on all sites $[k+1,(2m+1)k]$, where every $2k$ consecutive outcomes are determine recovery channels according to $(m_{x,1},m_{z,1},m_{x,2},m_{z,2},\dots,m_{x,k},m_{z,k})$, where $\sigma^{z/x}_{f_{n}}$ is applied to $f_{n}$ if $\sum_{s=1}^{m}m_{x/z,s}\cong 1$ mod $2$.
@@ 7 Outlook
We have shown that all “standard” teleportation protocols that teleport $k$ logical states a distance $L>vT$ (13) are equivalent to preparing a resource state with nontrivial SPT order protected by the symmetry $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ (65) using local unitary evolution for time $T\ll L$, measuring the $2k$ string order parameters (63) in the bulk, and applying Pauli recovery gates (34) based on the outcomes. These protocols are physical in that the logical states $\lvert\psi^{\,}_{n}\rangle$ are realized on particular physical qubits. We require that no unitaries are applied to a site after it is measured, though we expect that this requirement can be relaxed.
@@
The crucial feature of standard teleportation protocols, from the perspective of constraining the resource state, is that all outcome-dependent operations are Pauli gates and all classical processing is linear. The recovery group—generated by all physical error-correction operators $\chi$ in $\mathcal{R}$ (18), modulo complex phases—is simply the $k$-qubit Pauli group $\mathbf{P}^{\vphantom{*}}_{*}{k}/\mathsf{U}\left(1\right)$ for standard teleportation, which forms a projective representation of the symmetry group $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ (65). This simple structure of the recovery group ensures that the resource state $\lvert\Psi_{t}\rangle$ (36) belongs to a single class of SPT phases protected by the symmetry $\mathcal{G}=\mathbb{Z}_{2}^{2k}$. We are confident that our analysis can be extended, mutatis mutandis, to more general recovery operations, where the choice of recovery group dictates both the protecting symmetry $\mathcal{G}$ and phase of matter of the resource state, while the nature of the classical postprocessing of measurement outcomes dictates how that symmetry is represented on the physical qubits.
@@
In this sense, our results (namely, Theorem 18) represent a first step toward classifying all quantum teleportation protocols based solely on their recovery (or decoding) channels (18). There are numerous recovery operations to consider beyond the “Paulis and linear processing” of standard teleportation: The corresponding group can be continuous and non-Abelian (i.e., $\chi^{\,}_{g}\,\chi^{\,}_{h}\neq\omega(g,h)\,\chi^{\,}_{h}\,\chi^{\,}_{g}$ for some complex phase $\omega$), one can condition measurements (or the basis of a measurement) on prior outcomes, and one may condition operations on the outcomes of measurements in other ways. For example, Ref. 75 considers Pauli recovery operations, but requires nonlinear classical processing, in which case the resource state can be chosen to have SPT order, but the symmetry generators $U(g)$ are no longer “on site” (49). Other teleportation schemes use more complicated decoders to determine the appropriate recovery operations in the presence of, e.g., noisy measurement outcomes [12, 76, 30]. Finally, we note that matrix product states [18, 22] without a clear notion of string order (53[5] may be used as resource states for teleportation using a nondeterministic “repeat-until-success” recovery scheme. In general, we expect our approach using the Stinespring picture [42, 46] to be crucial to analyzing these more exotic protocols and their potential connection to SPT orders.
Lastly, we comment on the robustness of the teleportation schemes that fall under our formalism to noise and errors. It is known that teleportation protocols using SPT states are resources are robust to symmetry-respecting perturbations to the resource state [23, 27]. Therefore, by unveiling the presence of SPT order in all standard teleportation protocols, we have shown that all such protocols are robust to symmetry-respecting perturbations. Conversely, in the protocols we consider, precise knowledge of the measurement outcomes is crucial to the final recovery operations—there is no tolerance for even a single measurement error or generic error in the resource state. To achieve robustness to generic perturbations as well as noise and measurement errors, we need to consider the broader problem of fault-tolerant teleportation. There exist fault-tolerant teleportation protocols using 3D cluster states as resource states [16]. These have furthermore been connected to the presence of SPT order protected by higher-form symmetries [77]. From the perspective of our results, the ability to teleport using the 3D cluster state in the absence of errors follows from the presence of the membrane order parameters defined in Ref. 77, which play the role of string-order parameters. However, the fault tolerance of the protocol, and its relation to higher-form symmetries, lies beyond our framework, and is a compelling direction for future work.
@@ Acknowledgments
We thank Ollie Hart, Andy Lucas, Charles Stahl, and Chao Yin for useful discussions and collaboration on related work. This work was supported in part by the U.S. Air Force Office of Scientific Research under Grant No. FA9550-21-1-0195 (YH), the Office of Naval Research via Grant No. N00014-23-1-2533 (YH), the Simons Collaboration on Ultra-Quantum Matter via Award No. 651440 (DTS), and the US Department of Energy via Grant No. DE-SC0024324 (AJF).
@@ Appendix A Proof of Proposition 1
For the teleportation conditions (8) to be necessary and sufficient for quantum state transfer (see Def. 1), they must be equivalent to Eq. 5 for the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6) and final state $\lvert\Psi^{\,}_{T}\rangle$ (7). We prove both directions of implication in turn, using the fact that all operators acting on a qubit can be decomposed onto the Pauli operators $\mathds{1},X^{\vphantom{*}},Y^{\vphantom{*}},Z^{\vphantom{*}}$, which are spanned by $X^{\vphantom{*}}$ and $Z^{\vphantom{*}}$. The coefficients $\alpha$ and $\beta$ of a logical state $\lvert\psi\rangle$ can be extracted from the expectation values of the Pauli operators (4).
@@
We first establish that Eqs. 5, 6, and 7 imply Eq. 8—i.e., Eq. 8 is necessary for state transfer. First, Eq. 7 implies that $X^{\vphantom{*}}_{f_{n}}$ is the $X^{\vphantom{*}}$-type logical operator for the $n$th logical qubit acting on the final state $\lvert\Psi^{\,}_{T}\rangle$ (7), as it interchanges $\alpha^{\,}_{n}$ and $\beta^{\,}_{n}$. In the Heisenberg picture, this logical operator evolves under $\mathcal{W}$ according to
$X^{(n)}_{{\rm L}}(T)\,=\,X^{\vphantom{*}}_{f_{n}}(T)\,=\,\mathcal{W}^{\dagger}\,X^{\vphantom{*}}_{f_{n}}\,\mathcal{W}\,,~{}~{}$ (A.87)
@@
and now, using the fact that $X^{2}_{i_{n}}=\mathds{1}$ is involutory, we introduce an arbitrary operator $\mathsf{S}^{\,}_{n,x}$ via
$X^{(n)}_{{\rm L}}(T)\,$ = $=\,\mathds{1}\,\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f_{n}}\mathcal{W}\,=\,X^{2}_{i_{n}}\,\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f_{n}}\mathcal{W}$
$=\,X^{\vphantom{*}}_{i_{n}}\,\left(X^{\vphantom{*}}_{i_{n}}\,\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f_{n}}\mathcal{W}\right)\,\equiv\,X^{\vphantom{*}}_{i_{n}}\,\mathsf{S}^{\,}_{n,x}\,,~{}~{}$ (A.88)
which implicitly defines the (unitary) operator $\mathsf{S}^{\,}_{n,x}$ as
$\mathsf{S}^{\,}_{n,x}\,\equiv\,X^{\vphantom{*}}_{i_{n}}\,\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f_{n}}\mathcal{W}\,.$ (A.89)
@@
Leaving the coefficients $\alpha^{\,}_{j},\beta^{\,}_{j}$ for $j\neq n$ implicit for convenience, we apply $\mathsf{S}^{\,}_{n,x}$ (A.89) to the initial state $\lvert\Psi^{\,}_{0}(\alpha^{\,}_{n},\beta^{\,}_{n})\rangle$ (6) and use the first part of Eq. 5 to find
@@
$\mathsf{S}^{\,}_{n,x}\,\lvert\Psi^{\vphantom{\prime}}_{0}(\alpha^{\,}_{n},\beta^{\,}_{n})\rangle\,$ $=\,X^{\vphantom{*}}_{i_{n}}\,\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f_{n}}\mathcal{W}\,\lvert\Psi^{\vphantom{\prime}}_{0}(\alpha^{\,}_{n},\beta^{\,}_{n})\rangle$
@@
$=\,X^{\vphantom{*}}_{i_{n}}\,\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f_{n}}\lvert\Psi^{\vphantom{\prime}}_{T}(\alpha^{\,}_{n},\beta^{\,}_{n})\rangle$
@@
$=\,X^{\vphantom{*}}_{i_{n}}\,\mathcal{W}^{\dagger}\lvert\Psi^{\vphantom{\prime}}_{T}(\beta^{\,}_{n},\alpha^{\,}_{n})\rangle$
$=\,X^{\vphantom{*}}_{i_{n}}\lvert\Psi^{\vphantom{\prime}}_{0}(\beta^{\,}_{n},\alpha^{\,}_{n})\rangle$
$=\,\lvert\Psi^{\vphantom{\prime}}_{0}(\alpha^{\,}_{n},\beta^{\,}_{n})\rangle\,,~{}~{}$ (A.90)
@@
meaning that $\mathsf{S}^{\,}_{n,x}$ (A.89) is an element of the initial-state stabilizer group $\operatorname*{stab}\left(*\right){\lvert\Psi^{\,}_{0}\rangle}$ (3), and is guaranteed to be unitary. Next, applying the Heisenberg-evolved operator $X^{\vphantom{*}}_{f_{n}}(T)$ (A.87) to the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6) gives
$X^{(n)}_{{\rm L}}(T)\,\lvert\Psi^{\vphantom{\prime}}_{0}(\alpha^{\,}_{n},\beta^{\,}_{n})\rangle\,$ $=\,X^{\vphantom{*}}_{i_{n}}\,\mathsf{S}^{\,}_{n,x}\,\lvert\Psi^{\vphantom{\prime}}_{0}(\alpha^{\,}_{n},\beta^{\,}_{n})\rangle$
$=\,X^{\vphantom{*}}_{i_{n}}\,\lvert\Psi^{\vphantom{\prime}}_{0}(\alpha^{\,}_{n},\beta^{\,}_{n})\rangle\,,~{}~{}$ (A.91)
@@
which holds for any choice of $\alpha^{\,}_{n},\beta^{\,}_{n}$. This implies that $X^{(n)}_{{\rm L}}(T)\,=\,\mathcal{W}^{\dagger}\,X^{\vphantom{*}}_{f_{n}}\,\mathcal{W}\,$ realizes an $X^{\vphantom{*}}$-type logical operator for the $n$th logical qubit acting on the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6)—i.e., it acts as $X^{\vphantom{*}}_{i_{n}}$ times some stabilizer $\mathsf{S}^{\,}_{n,x}$ (A.89), where the latter acts trivially on $\lvert\Psi^{\,}_{0}\rangle$ (6).